# Observing hidden neuronal states in experiments

### v2

@@ Abstract
Tracking non-observable (hidden) states and their stability boundaries in experimental settings bridges the gap between real-world phenomena and nonlinear science. Specifically, closed-loop control methods with slow variations of feedback reference signals enable to dissect the underlying states of multi-scale complex systems. Following this strategy, we demonstrate the feasibility of tracking a family of neuronal steady states (stable and unstable) via variations of reference signal and reparameterizing the obtained curve by feedback current. Crucially, the method is versatile in that it applies to different cell types, both pyramidal (PY) neurons (like in Fig. 1), of excitability type I or II (Figs. 3(a) and 4), and inhibitory interneurons (Fig. 3(b)).
@@
These results can be extended to more complex neuronal states (e.g., limit cycles) by combining two recent advances, namely: 1. dynamic-clamp electrophysiology [33, 34, 35, 2], which allows for a two-way real-time communication between a neuronal tissue (e.g., neuron, network or neuronal sub-processes) and a computer simulation of a component of the neuronal tissue. This introduces the full range of feedback control theory and engineering techniques into electrophysiology; 2. Control-based continuation for experiments (CBCE, which combines feedback control theory and pseudo-arclength numerical continuation for tracking solution branches of nonlinear systems directly from noisy experimental data [7]. Indeed, CBCE method has been successfully applied in various experimental systems, for instance, in mechanical vibration and buckling experiments [8], pedestrian flow experiments [9], cylindrical pipe flow simulations [11] and feasibility studies for synthetic gene networks [12, 13]. Noteworthy, the methodology presented in this work can be seen as a special case of CBCE. Future work will focus on employing the full CBCE to track more complex neuronal states. This is crucial as it will allow to validate computational models by comparing their numerical bifurcation diagrams with experimental ones. This approach will have direct impact in experimental labs enabling biologists to have access to novel states with which to carry out new experimental paradigms. We envisage this extended protocol could be used to develop next-generation closed-loop deep-brain stimulation devices to treat certain brain diseases such as epilepsy [36], where it could help monitor control the excitability threshold of key neural populations in real time.
Materials and Method.–
Protocols.–
@@
Male Wistar rats were used in this study (age P21, N=8 animals). For the bifurcation diagrams (see Fig. 1(b) and Fig. SI1), 4 of these 8 rats were used. PY Cells 1-3 were recorded from different animals, PY cells 4-5 were recorded from the same animal. To study the effects of hysteresis (see Fig. 2) 20 neurons were recorded from the 4 remaining animals (221222, 230110, 230111, 230112). The control group were 6 neurons from 2 rats (230110, 230112). The QX group were 8 neurons from 4 rats (221222, 230110, 230111, 230112). The QX+Cd group were 6 neurons from 4 rats (221222, 230110, 230111, 230112).
@@
The use and handling of animals was performed in accordance with the European Community Council Directive 86/609/EEC. Horizontal 300-µm-thick brain slices were prepared as described in our previous studies [37]. The slices contained the hippocampus and the adjacent cortical regions and were kept in the artificial cerebrospinal fluid (ACSF) with the following composition (in mM): 126 NaCl, 24 NaHCO3, 2.5 KCl, 2 CaCl2, 1.25 NaH2PO4, 1 MgSO4, 10 glucose (bubbled with 95% O2/5% CO2 gas mixture). All the listed chemicals were purchased from Sigma-Aldrich (St. Louis, MO, USA).
@@
We applied first VC and then CC protocol to $5$ neurons in the entorhinal cortex of $4$ male Wistar rats [37]. We first performed the VC neuronal recordings varying the hold voltage $V_{\mathrm{h}}$ from $-80$ mV to $+30$ mV slowly with $\dot{V}_{\mathrm{h}}=1.83$mV///s, while measuring current, called $I_{\mathrm{vc}}$ in Fig. 1(b,c). Subsequently, for the CC recordings we first determine the minimal injected current required to induce a depolarization block of action potential generation ($I_{\mathrm{db}}$ in Fig. 1(b), upper limit of current input where firing occurs). Then we injected hold current $I_{\mathrm{h}}$, gradually increasing from 00 pA to $I_{\mathrm{db}}$ during $60$ s, such that $\dot{I}_{\mathrm{h}}$ is in the range $6.6\ldots 7.4$ pA///s for the $5$ neurons, while recording voltage $V_{\mathrm{cc}}(t)$. See SI for further checks (e.g., for hysteresis).
Computational models.–
We applied the VC and the CC protocols, with slow variations in the feedback reference signal and in the applied current, respectively, to the classical neuron model adapted to many contexts, namely the Morris-Lecar model [30] (ML), and to a standard neuron model that was specifically designed to account for inhibitory neural activity, namely the Wang-Buzsaki model [31] (WB).
@@
To obtain Figs. 3(a) and 4, we have used the following parameter values.
@@
value type-I 20 2 -60 4.0 120 12 -84 40 0.067 -1.2 18 12 17.4
value type-II 20 2 -60 4.4 120 12 -84 150 0.040 -1.2 18 2 30.0
@@
To obtain Fig. 3(b) of the main manuscript, we have used the following parameter values.
Table 2: Parameter values for the Wang-Buzsaki model (10).
@@
Data and code availability.–
All presented data, as well as, codes to process them, are available via the following Github repository: https://github.com/serafimrodrigues/Amakhin-etal-2025/.
@@
Supplementary Information (SI)
S1 Experimental bifurcation diagrams for Pyramidal (PY) cells 1–4, including current and voltage clamp ramp data
Figure S1: Experimental bifurcation diagrams for PY cells 1–4. Protocols identical to Fig. 1(b) in main manuscript.
Figure S1 presents the recordings for VC and CC protocol for PY cells 1–4 in the same way as Fig. 1(b) in the main manuscript (which was for PY cell 5). The graphs superimpose the observations of our VC protocol with slow variation of the feedback reference signal and the CC protocol with slow variation of the externally applied currents. The graphs confirm (as does Fig. 1(c) from the main manuscript) that type-II properties of the neurons vary between individual cells. We also observe that in all PY cells for $I_{\mathrm{h}}>I_{\mathrm{db}}$ and for PY cell 3 for small $I_{\mathrm{h}}$ the distance between the stable steady states obtained from the VC protocol and the corresponding trajectory segments obtained from the CC protocol is larger than the bias caused by dynamic variation of $V_{\mathrm{h}}$ and $I_{\mathrm{h}}$ respectively, based on estimate (5) of the main text. We do not have a complete physiological explanation for this mismatch, but we hypothesise that this is due to the natural physiological drift of underlying neuronal processes as one performs successive recordings on the same PY cell (see also Fig. 2 of the main manuscript for a comparison between successive ramp-up and ramp-down on the same PY cell).
@@
A characteristic for the class of excitability is the dependence of the spiking frequency on the external current $I_{\mathrm{ext}}$ as $I_{\mathrm{ext}}$ approaches the lower limit of the spiking region, with a drop in frequency expected for type-I neurons. Fig. S2 presents the interspike interval as a function of the hold current $I_{\mathrm{h}}$ in panel (a), and the frequency against $I_{\mathrm{h}}$ in panel (b) for the 5 PY cells for which experimental bifurcation diagrams are shown in Fig. 1 of the main manuscript and Fig. S1. We detected crossing times for $V_{\mathrm{cc}}(t)$ through the threshold $V=0$ from below (or equivalently Poincaré map, at $V=0$). As visible in Fig. S2, even though the 5 PY cells have different steady-state curves (obtained with the VC protocol), their frequency behaviours during the CC protocol are qualitatively similar with drops from $\approx 15$ Hz to $\lesssim 5$ Hz for all PY cells. Hence, the frequency drop does not allow us to conclusively distinguish between different excitability classes for the present PY cells. However, the steady-state bifurcation diagrams obtained through VC protocols show that PY cells 3-5 exhibit an experimental bifurcation diagram that is compatible with type-1 excitability, while PY cells 1 and 2 have type-2 compatible steady-state curves.
Figure S2: Interspike intervals in seconds (panel (a)) and frequency in Hz (panel (b)) for the spiking responses during the CC protocols with slowly-varying applied current near the low-current limit of the spiking region.
@@
When performing the VC protocol with slow variation of the feedback reference signal we observe fluctuations around the states that would be unstable in the CC protocol. The highlighted ellipse in Fig. S3(a) shows a prominent example of large-scale fluctuations. We observe this phenomenon in all PY cells. Figure S3(b) shows details of the typical time profile of these fluctuations in the measured feedback current $I_{\mathrm{vc}}=g_{c}(V_{h}-V)$ in blue. In the background Fig. S3(b) shows a neuronal firing pattern (in orange) around the same currents to give a visual comparison of fluctuation amplitude and frequency with the natural neuronal firing patterns. We also compute the moving average of $I_{\mathrm{vc}}(t)$ (dashed blue), which shows the equilibrium level of $I_{\mathrm{vc}}$. We observe that between the large fluctuations the feedback current $I_{\mathrm{vc}}$ stabilises to this equilibrium level with a consistent rate of decay and consistently smaller levels of smaller fluctuations. We used the observed decay rate toward the mean to estimate the decay half-time $\tau_{1/2}$ in eq. (5) of the main manuscript.
To simulate the intermittent fluctuations around unstable steady states observed in the experimental recordings (see Fig. S3), we used the Morris-Lecar model (eq. (8) of the main manuscript) with added external current spikes in the $V$-equation. Namely, we replaced the $V$-equation from eq. (8) by
@@
where $I_{i}(t)=I_{p}\delta(t-t_{i})$ is a current spike modelled by a Dirac delta input, which we approximated by two Heaviside functions. Namely: $\delta(t-t_{i})\approx\text{Heav}(t-t_{i})\ast\text{Heav}(t_{i}+t_{\mathrm{dur}}-t)$, with the duration of the spike chosen to be $t_{\mathrm{dur}}=10$ms, and its amplitude being $I_{p}=50$pA. We also added noise in the form of a Wiener process $W$ of amplitude $\sigma_{W}=0.5$. Other parameters are as in Table 1. For the loss of contact simulation (Fig. S4(b)) we used a piecewise-constant $g_{c}$ dropping from $20$ to $5$ along the unstable branch of equilibria (other parameters as in Table 1) and we also added noise using a Wiener process $W$ as in the previous case, of amplitude $\sigma_{W}=0.1$.
Figure S3: Experimental intermittent fluctuations. (a) Reproduction of Fig. 1(b) of main manuscript with the region of intermittent fluctuations highlighted by a black ellipse. (b) Fluctuation of the current $I_{\mathrm{vc}}$ from the VC protocol, with apparent intermittent fluctuation (blue), compared with the voltage spikes obtained from the CC protocol (orange). Red dashed lines illustrate how we compute $\tau_{1/2}$.
Figure S4: Reproducing the experimental intermittent fluctuations with the ML model. (a-b) Simulation of the ML model with current spikes artificially added to the voltage equation, viewed in the $(I_{\mathrm{vc}},V_{\mathrm{h}})$-plane in panel (a) and with the time series of $I_{\mathrm{vc}}$ in panel (b). (c) Simulation of the ML model with artificial loss of contact through a drop of gain $g_{\mathrm{c}}$, viewed in the $(I_{\mathrm{vc}},V_{\mathrm{h}})$-plane.
@@
This can be understood by first comparing I-V relationships of voltage-gated channels of neurons in slice preparations versus isolated neurons, and subsequently inferring what occurs to isolated voltage-gated channels that are poorly controlled due to space-clamp error. Neurons in slice preparations have lower input resistance and higher membrane capacitance when compared to an isolated neuron [38]. We observed the same effect with our entorhinal cortical neurons, which in slice preparations had $R\in\left]150,350\right[M\Omega$ and $C\approx 100$ pF, while in a previous study (unpublished) the isolated case had $R\in\left]500,1000\right[M\Omega$ and $C\in\left]10,30\right[$ pF. Thus, in slice preparations the typically observed voltage range in which the I-V relationship has a negative slope (e.g., for sodium channels) is quite large [39]. In ideal conditions (i.e., when voltage-gated channels are more or less clustered within a neuronal site, for instance, close to the soma), space-clamp error is minimised. In this case, when the membrane voltage is slowly increased then individual negative peak currents of voltage-gated channels are successively activated (i.e., a kind of averaging effect occurs due to the clustering of channels), which give rise to a wide voltage range in which the I-V relationship has a negative slope. In other words, the current peaks are distributed across the voltage range of the I-V relationship [39].
@@
Noteworthy, the fast sodium current peaks occur during this negative slope phase because the voltage dependence of the transient sodium channel activation (which mostly mediate these peaks) is similar to the activation of the persistent sodium channels. The poorly voltage-clamped transient sodium channels mediate the fast negative currents, while the transient ones mediate a slow negative current, both of which appear in the same voltage range [40]. In contrast, isolated neurons typically lose most of their thin dendritic branches (i.e., abbreviated morphology), hence voltage-gated channels (e.g., sodium current) display only one negative peak threshold current under a step depolarisation [38, 41]. From these observations, it is reasonable to assume that in non-ideal conditions (i.e., space-clamp error) there are isolated channels furthest from the cell soma or electrode contact (e.g., distal dendrites with thin branches, or dendritic spines with increased resistance of the cytoplasm, or distal sodium channels) [42, 43]. These isolated channels have I-V relationships with negative slope but with a shorter voltage range and distinct activation threshold. Thus, under a VC protocol the mismatch between the hold voltage ($V_{\mathrm{h}}$) and the real-membrane voltage ($V$) of distal voltage-gated channels can be very large, resulting in the observed intermittent fluctuations. Additionally to the above mentioned factors, the delay of the feedback (associated with the amplifier, which is in the order tens of nanoseconds) is likely to play a role in these observed intermittent fluctuations. However, from a theoretical standpoint, a fixed delay can induce a Hopf bifurcation, and therefore, robust oscillations. So instead, we envisage a varying delay, due to possibly latent processes, could create this effect. Of course, to really address this point, further experiments are required.
@@
To simulate the above hypothesis we added additional instantaneous currents (i.e., to model the effect of non-ideally clamped channels as external forcing that correspond to non-homogeneous terms); see Section Computation models of the main manuscript for details, in particular eq. (10). The results are depicted in Fig. S4 in the $(I,V)$-plane (panel (a) and as a corresponding time series $I_{\mathrm{vc}}(t)$ (panel (b)). Both, time profile and $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$-curve closely resemble the fluctuations in the experimental data in Fig. S3(a). Although, these simulations may explain the experimental observations, we do not exclude other possible mechanisms (or a mix of mechanisms). To this end, note that the standard VC protocol was designed to stabilise the membrane voltage $V$ at some static reference signal $V_{h}$, for a given, large enough, control gain $g_{c}$. However, our VC protocol involves slowly varying $V_{h}$, which causes the neuronal membrane $V$ to quasi-statically stabilise (drift) without ever reaching the true steady state. Along the unstable branch, the repulsion (associated with unstable eigenvalues) may not necessarily be constant and thus the degree of effectiveness of the control signal $g_{c}(V_{h}-V)$ may vary. Noteworthy, this control signal is effectively the proportional control term in the proportional–integral–derivative (PID) control formalism. Thus, it is also possible the proportional control does not achieve a full control of more complex physiological processes in which case the control may only be intermittent. Other possibilities may include temporary loss of contact between the electrode and the neuronal membrane. To elucidate these points, we simulated equation (8) but with variations in $g_{c}$. Specifically, we considered a piecewise-constant $g_{c}$ corresponding to decreasing its default value (thus modelling loss of control) along the unstable branch of equilibrium; see Fig. S4(c) for a simulation. While this scenario bears less resemblance to the experimental observation, it can either be used to rule out this possibility or to consider new experiments to further clarify these observations.
@@
The voltage-clamp recordings of the responses to the slow ramps of command voltage (Fig. S5 (a)) were made to obtain the current-voltage relationships (I-V relationships). With the potassium gluconate-based pipette solution being used, the I-V relationship obtained by applying the ascending ramp was not the same as the one obtained by applying the descending ramp (Fig. S5 (b)). The ascending ramp produced an S-shaped I-V relationship, which had the negative slope in the voltage range of $-50$ to $-30$ mV (Fig. S5 (b), CTRL, green trace). The descending ramp of the command voltage produced an I-V curve with a positive slope throughout the physiological voltage range.

### v3

@@ 3 Discussion
Tracking non-observable (hidden) states and their stability boundaries in experimental settings bridges the gap between real-world phenomena and nonlinear science. Specifically, closed-loop control methods with slow variations of feedback reference signals enable to dissect the underlying states of multi-scale complex systems. Following this strategy, we demonstrate the feasibility of tracking a family of neuronal steady states (stable and unstable) via variations of reference signal and reparameterizing the obtained curve by feedback current. The method is versatile in that it applies to different cell types, both pyramidal (PY) neurons (like in Fig 1), of excitability class I or II (Figs 3A and 4), and inhibitory interneurons (Fig 3B).
Our analysis and the resulting estimate for the nonzero-speed induced bias during the VC protocol in (5) indicates that for the steady states detected in VC ramps our ramp speed (less than 33 mV/\!/s) does not create a bias that is notable when compared to drift of cell properties over time or fluctuations caused by external effects (such as the dropouts visible in Fig 1B). For the CC protocol and parameter areas were periodic spiking is stable, transient analysis for step current responses results in estimates such as (9), where bias caused by nonzero current ramp speed ($\approx 7.5$ pA//s) is also negligible. This implies that faster ramp speeds are permissible when optimising trade-off between drift of cell properties and bias due to nonzero ramp speed. The single-ramp nature of the CC protocol leaves the question of stability open for some parts of the steady-state branch. These results can be improved by varying the hold current $I_{\mathrm{h}}$ non-monotonically and be switching from VC to CC protocol at selected parts of the steady-state branch determined by the VC protocol.
@@
These results can be also extended to more complex neuronal states (e.g. limit cycles) by combining two recent advances, namely: 1. dynamic-clamp electrophysiology[sharp1993, marder1996, chizhov2014, kirst2015, ori2020, pfeiffer2022], which allows for a two-way real-time communication between a neuronal tissue (e.g. neuron, network or neuronal sub-processes) and a computer simulation of a component of the neuronal tissue. This introduces the full range of feedback control theory and engineering techniques into electrophysiology; 2. Control-based continuation for experiments (CBCE, which combines feedback control theory and pseudo-arclength numerical continuation for tracking solution branches of nonlinear systems directly from noisy experimental data [sieber2008a]. Indeed, CBCE method has been successfully applied in various experimental systems, for instance, in mechanical vibration and buckling experiments [RENSON2019449], pedestrian flow experiments [panagiotopoulos2022control], cylindrical pipe flow simulations [willis_duguet_omelchenko_wolfrum_2017] and feasibility studies for synthetic gene networks [de2022control, blyth2023numerical]. Noteworthy, the methodology presented in this work can be seen as a special case of CBCE. Future work will focus on employing the full CBCE to track more complex neuronal states, for instance, unstable spiking states in single neurons or unstable traveling-wave states in neural tissue using micro-electrode arrays (MEA). This is crucial as it will allow us to validate computational models by comparing their numerical bifurcation diagrams with experimental ones; it will also help obtaining better model fitting based on both voltage-clamp and current-clamp data. This approach will have direct impact in experimental labs enabling biologists to have access to novel states with which to carry out new experimental paradigms. We envisage this extended protocol could be used to develop next-generation closed-loop deep-brain stimulation devices to treat certain brain diseases such as epilepsy [breakspear2006], where it could help monitor control the excitability threshold of key neural populations in real time.
4 Materials and Methods
4.1 Ethics Statement
All animal treatments were authorized by the Sechenov Institute of Evolutionary Physiology and Biochemistry Bioethics Committee (protocol no. 1-7/2022, 27 January 2022) and adhered to the European Community Council Directive 86/609/EEC.
4.2 Protocols
@@
Male Wistar rats were used in this study (age P21, N=8 animals). For the bifurcation diagrams (see Figs 1B and A in S1 Text), 4 of these 8 rats were used. PY Cells 1-3 were recorded from different animals, PY cells 4-5 were recorded from the same animal. To study the effects of hysteresis (see Fig 2) 20 neurons were recorded from the 4 remaining animals (221222, 230110, 230111, 230112). The control group were 6 neurons from 2 rats (230110, 230112). The QX group were 8 neurons from 4 rats (221222, 230110, 230111, 230112). The QX+Cd group were 6 neurons from 4 rats (221222, 230110, 230111, 230112).
@@
The use and handling of animals was performed in accordance with the European Community Council Directive 86/609/EEC. Horizontal 300-µm-thick brain slices were prepared as described in our previous studies [amakhin2021]. The slices contained the hippocampus and the adjacent cortical regions and were kept in the artificial cerebrospinal fluid (ACSF) with the following composition (in mM): 126 NaCl, 24 NaHCO3, 2.5 KCl, 2 CaCl2, 1.25 NaH2PO4, 1 MgSO4, 10 glucose (bubbled with 95% O2/5% CO2 gas mixture). All the listed chemicals were purchased from Sigma-Aldrich (St. Louis, MO, USA).
@@
We applied first VC and then CC protocol to $5$ neurons in the entorhinal cortex of $4$ male Wistar rats [amakhin2021]. We first performed the VC neuronal recordings varying the hold voltage $V_{\mathrm{h}}$ from 80-80 mV to +30+30 mV slowly with $\dot{V}_{\mathrm{h}}=1.83$ mV//s, while measuring current, called $I_{\mathrm{vc}}$ in Figs 1B and 5C. Subsequently, for the CC recordings we first determine the minimal injected current required to induce a depolarization block of action potential generation ($I_{\mathrm{db}}$ in Fig 1B, upper limit of current input where firing occurs). Then we injected hold current $I_{\mathrm{h}}$, gradually increasing from 0 pA to $I_{\mathrm{db}}$ during $60$ s, such that $\dot{I}_{\mathrm{h}}$ is in the range $6.6\ldots 7.4$ pA//s for the $5$ neurons, while recording voltage $V_{\mathrm{cc}}(t)$. See S1 Text for further checks (e.g., for hysteresis).
4.3 Computational models
We applied the VC and the CC protocols, with slow variations in the feedback reference signal and in the applied current, respectively, to the classical neuron model adapted to many contexts, namely the Morris-Lecar model [morris1981] (ML), and to a standard neuron model that was specifically designed to account for inhibitory neural activity, namely the Wang-Buzsáki model [wang1996] (WB). Parameter values are given in Tables 1 and 2 below, respectively.
@@ 4.3 Computational models
To obtain Figs 3A and 4, we have used the following parameter values.
@@
value class-I 20 2 -60 4.0 120 12 -84 40 0.067 -1.2 18 12 17.4
value class-II 20 2 -60 4.4 120 12 -84 150 0.040 -1.2 18 2 30.0
@@
To obtain Fig 3B, we have used the following parameter values.
Table 2: Parameter values for the Wang-Buzsáki model (12).
@@
Data and code availability
All presented data, as well as, codes to process them, are available via the following Github repository:
https://github.com/serafimrodrigues/Amakhin-etal-2025/.
@@ Acknowledgments
S1 Text
S1 Experimental bifurcation diagrams for Pyramidal (PY) cells 1–5, including current and voltage clamp ramp data
Figure A in S1 Text: Experimental bifurcation diagrams for PY cells 1–5 (panel E identical to Fig 1B of main manuscript). Protocols and color codes identical to Fig 1B in main manuscript.
Fig A in S1 Text presents the recordings for VC and CC protocol for all PY cells (1–5) in the same way as Fig 1B in the main manuscript (which was only for PY cell 5). The graphs superimpose the observations of our VC protocol with slow variation of the feedback reference signal and the CC protocol with slow variation of the externally applied currents. The graphs confirm (as does Fig 1C from the main manuscript) that (Hodgkin excitability) class-II properties of the neurons vary between individual cells. We also observe that in all PY cells for $I_{\mathrm{h}}>I_{\mathrm{db}}$ and for PY cell 3 for small $I_{\mathrm{h}}$ the distance between the stable steady states obtained from the VC protocol and the corresponding trajectory segments obtained from the CC protocol is larger than the bias caused by dynamic variation of $V_{\mathrm{h}}$ and $I_{\mathrm{h}}$ respectively, based on estimate (5) of the main text. We do not have a complete physiological explanation for this mismatch, but we hypothesise that this is due to the natural physiological drift of underlying neuronal processes as one performs successive recordings on the same PY cell (see also Fig 2 of the main manuscript for a comparison between successive ramp-up and ramp-down on the same PY cell).
@@
A characteristic for the class of excitability is the dependence of the spiking frequency on the external current $I_{\mathrm{ext}}$ as $I_{\mathrm{ext}}$ approaches the lower limit of the spiking region, with a drop in frequency expected for class-I neurons. Fig B in S1 Text presents the interspike interval as a function of the hold current $I_{\mathrm{h}}$ in panel A, and the frequency against $I_{\mathrm{h}}$ in panel B for the 5 PY cells for which experimental bifurcation diagrams are shown in Fig 1 of the main manuscript and Fig A in S1 Text. We detected crossing times for $V_{\mathrm{cc}}(t)$ through the threshold $V=0$ from below (or equivalently Poincaré map, at $V=0$). As visible in Fig B in S1 Text, even though the 5 PY cells have different steady-state curves (obtained with the VC protocol), their frequency behaviours during the CC protocol are qualitatively similar with drops from $\approx 15$ Hz to $\lesssim 5$ Hz for all PY cells. Hence, the frequency drop does not allow us to conclusively distinguish between different excitability classes for the present PY cells. However, the steady-state bifurcation diagrams obtained through VC protocols show that PY cells 3-5 exhibit an experimental bifurcation diagram that is compatible with class-1 excitability, while PY cells 1 and 2 have class-2 compatible steady-state curves.
Figure B in S1 Text: Interspike intervals in seconds (panel A) and frequency in Hz (panel B) for the spiking responses during the CC protocols with slowly-varying applied current near the low-current limit of the spiking region.
@@
When performing the VC protocol with slow variation of the feedback reference signal we observe fluctuations around the states that would be unstable in the CC protocol. The highlighted ellipse in Fig CA in S1 Text shows a prominent example of large-scale fluctuations. We observe this phenomenon in all PY cells. Fig CB in S1 Text shows details of the typical time profile of these fluctuations in the measured feedback current $I_{\mathrm{vc}}=g_{c}(V_{h}-V)$ in blue. In the background Fig CB in S1 Text shows a neuronal firing pattern (in orange) around the same currents to give a visual comparison of fluctuation amplitude and frequency with the natural neuronal firing patterns. We also compute the moving average of $I_{\mathrm{vc}}(t)$ (dashed blue), which shows the equilibrium level of $I_{\mathrm{vc}}$. We observe that between the large fluctuations the feedback current $I_{\mathrm{vc}}$ stabilises to this equilibrium level with a consistent rate of decay and consistently smaller levels of smaller fluctuations. We used the observed decay rate toward the mean to estimate the decay half-time $\tau_{1/2}$ in eq. (5) of the main manuscript.
To simulate the intermittent fluctuations around unstable steady states observed in the experimental recordings (see Fig C in S1 Text), we used the Morris-Lecar model (eq (8) of the main manuscript) with added external current spikes in the $V$-equation. Namely, we replaced the $V$-equation from eq (8) by
@@
where $I_{i}(t)=I_{p}\delta(t-t_{i})$ is a current spike modelled by a Dirac delta input, which we approximated by two Heaviside functions. Namely: $\delta(t-t_{i})\approx\text{Heav}(t-t_{i})\ast\text{Heav}(t_{i}+t_{\mathrm{dur}}-t)$, with the duration of the spike chosen to be $t_{\mathrm{dur}}=10$ms, and its amplitude being $I_{p}=50$pA. We also added noise in the form of a Wiener process $W$ of amplitude $\sigma_{W}=0.5$. Other parameters are as in Materials and Methods of the main manuscript. For the loss of contact simulation (Fig DB in S1 Text) we used a piecewise-constant $g_{c}$ dropping from $20$ to $5$ along the unstable branch of equilibria (other parameters as in Materials and Methods of the main manuscript) and we also added noise using a Wiener process $W$ as in the previous case, of amplitude $\sigma_{W}=0.1$.
Figure C in S1 Text: Experimental intermittent fluctuations. A: Reproduction of Fig 1B of main manuscript with the region of intermittent fluctuations highlighted by a black ellipse. B: Fluctuation of the current $I_{\mathrm{vc}}$ from the VC protocol, with apparent intermittent fluctuation (blue), compared with the voltage spikes obtained from the CC protocol (orange). Red dashed lines illustrate how we compute $\tau_{1/2}$.
Figure D in S1 Text: Reproducing the experimental intermittent fluctuations with the ML model. A-B: Simulation of the ML model with current spikes artificially added to the voltage equation, viewed in the $(I_{\mathrm{vc}},V_{\mathrm{h}})$-plane in panel A and with the time series of $I_{\mathrm{vc}}$ in panel B. C: Simulation of the ML model with artificial loss of contact through a drop of gain $g_{\mathrm{c}}$, viewed in the $(I_{\mathrm{vc}},V_{\mathrm{h}})$-plane.
@@
This can be understood by first comparing I-V relationships of voltage-gated channels of neurons in slice preparations versus isolated neurons, and subsequently inferring what occurs to isolated voltage-gated channels that are poorly controlled due to space-clamp error. Neurons in slice preparations have lower input resistance and higher membrane capacitance when compared to an isolated neuron [kay1986]. We observed the same effect with our entorhinal cortical neurons, which in slice preparations had $R\in\left]150,350\right[M\Omega$ and $C\approx 100$ pF, while in a previous study (unpublished) the isolated case had $R\in\left]500,1000\right[M\Omega$ and $C\in\left]10,30\right[$ pF. Thus, in slice preparations the typically observed voltage range in which the I-V relationship has a negative slope (e.g., for sodium channels) is quite large [sherman1999]. In ideal conditions (i.e., when voltage-gated channels are more or less clustered within a neuronal site, for instance, close to the soma), space-clamp error is minimised. In this case, when the membrane voltage is slowly increased then individual negative peak currents of voltage-gated channels are successively activated (i.e., a kind of averaging effect occurs due to the clustering of channels), which give rise to a wide voltage range in which the I-V relationship has a negative slope. In other words, the current peaks are distributed across the voltage range of the I-V relationship [sherman1999].
@@
Noteworthy, the fast sodium current peaks occur during this negative slope phase because the voltage dependence of the transient sodium channel activation (which mostly mediate these peaks) is similar to the activation of the persistent sodium channels. The poorly voltage-clamped transient sodium channels mediate the fast negative currents, while the transient ones mediate a slow negative current, both of which appear in the same voltage range [fleidervish1996]. In contrast, isolated neurons typically lose most of their thin dendritic branches (i.e., abbreviated morphology), hence voltage-gated channels (e.g., sodium current) display only one negative peak threshold current under a step depolarisation [kay1986, golding1999]. From these observations, it is reasonable to assume that in non-ideal conditions (i.e., space-clamp error) there are isolated channels furthest from the cell soma or electrode contact (e.g. distal dendrites with thin branches, or dendritic spines with increased resistance of the cytoplasm, or distal sodium channels) [dumenieu2017, petersen2015]. These isolated channels have I-V relationships with negative slope but with a shorter voltage range and distinct activation threshold. Thus, under a VC protocol the mismatch between the hold voltage ($V_{\mathrm{h}}$) and the real-membrane voltage ($V$) of distal voltage-gated channels can be very large, resulting in the observed intermittent fluctuations. Additionally to the above mentioned factors, the delay of the feedback (associated with the amplifier, which is in the order tens of nanoseconds) is likely to play a role in these observed intermittent fluctuations. However, from a theoretical standpoint, a fixed delay can induce a Hopf bifurcation, and therefore, robust oscillations. So instead, we envisage a varying delay, due to possibly latent processes, could create this effect. Of course, to really address this point, further experiments are required.
@@
To simulate the above hypothesis we added additional instantaneous currents (i.e. to model the effect of non-ideally clamped channels as external forcing that correspond to non-homogeneous terms); see Section Computation models of the main manuscript for details, in particular eq. (10). The results are depicted in Fig D in S1 Text in the $(I,V)$-plane (panel A) and as a corresponding time series $I_{\mathrm{vc}}(t)$ (panel B). Both, time profile and $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$-curve closely resemble the fluctuations in the experimental data in Fig CA in S1 Text. Although, these simulations may explain the experimental observations, we do not exclude other possible mechanisms (or a mix of mechanisms). To this end, note that the standard VC protocol was designed to stabilise the membrane voltage $V$ at some static reference signal $V_{h}$, for a given, large enough, control gain $g_{c}$. However, our VC protocol involves slowly varying $V_{h}$, which causes the neuronal membrane $V$ to quasi-statically stabilise (drift) without ever reaching the true steady state. Along the unstable branch, the repulsion (associated with unstable eigenvalues) may not necessarily be constant and thus the degree of effectiveness of the control signal $g_{c}(V_{h}-V)$ may vary. Noteworthy, this control signal is effectively the proportional control term in the proportional–integral–derivative (PID) control formalism. Thus, it is also possible the proportional control does not achieve a full control of more complex physiological processes in which case the control may only be intermittent. Other possibilities may include temporary loss of contact between the electrode and the neuronal membrane. To elucidate these points, we simulated the Morris-Levar equation (see Material and Methods in main manuscript) but with variations in $g_{c}$. Specifically, we considered a piecewise-constant $g_{c}$ corresponding to decreasing its default value (thus modelling loss of control) along the unstable branch of equilibrium; see Fig DC in S1 Text for a simulation. While this scenario bears less resemblance to the experimental observation, it can either be used to rule out this possibility or to consider new experiments to further clarify these observations.
@@
The voltage-clamp recordings of the responses to the slow ramps of command voltage (Fig EA in S1 Text) were made to obtain the current-voltage relationships (I-V relationships). With the potassium gluconate-based pipette solution being used, the I-V relationship obtained by applying the ascending ramp was not the same as the one obtained by applying the descending ramp (Fig EB in S1 Text). The ascending ramp produced an S-shaped I-V relationship, which had the negative slope in the voltage range of 50-50 to 30-30 mV (Fig EB in S1 Text, CTRL, green trace). The descending ramp of the command voltage produced an I-V curve with a positive slope throughout the physiological voltage range.